# Hypencoder: Hypernetworks for Information Retrieval

### v1

@@ Abstract.
The vast majority of retrieval models depend on vector inner products to produce a relevance score between a query and a document. This naturally limits the expressiveness of the relevance score that can be employed. We propose a new paradigm, instead of producing a vector to represent the query we produce a small neural network which acts as a learned relevance function. This small neural network takes in a representation of the document, in this paper we use a single vector, and produces a scalar relevance score. To produce the little neural network we use a hypernetwork, a network that produce the weights of other networks, as our query encoder or as we call it a Hypencoder. Experiments on in-domain search tasks show that Hypencoder is able to significantly outperform strong dense retrieval models and has higher metrics then reranking models and models an order of magnitude larger. Hypencoder is also shown to generalize well to out-of-domain search tasks. To assess the extent of Hypencoder’s capabilities, we evaluate on a set of hard retrieval tasks including tip-of-the-tongue retrieval and instruction-following retrieval tasks and find that the performance gap widens substantially compared to standard retrieval tasks. Furthermore, to demonstrate the practicality of our method we implement an approximate search algorithm and show that our model is able to search 8.8M documents in under 60ms.
journalyear: 2025copyright: rightsretained
@@ 1. Introduction
Efficient neural retrieval models are based on a bi-encoder (or two tower) architecture, in which queries and documents are represented separately using either high-dimensional sparse (Zamani et al., 2018; Formal et al., 2021, 2022) or relatively low-dimensional dense vectors (Karpukhin et al., 2020; Xiong et al., 2021; Khattab and Zaharia, 2020; Zeng et al., 2022; Gao and Callan, 2022). These models use simple and light-weight similarity functions, e.g., inner product or cosine similarity, to compute the relevance score for a given pair of query and document representations. We demonstrate theoretically that inner product similarity functions fundamentally limit the types of relevance that retrieval models can express. Specifically, we prove that there is always a set of relevant documents which cannot be perfectly retrieved regardless of the query vector and specific encoder model.
@@
Motivated by this theoretical argument, we introduce a new category of retrieval models that can capture complex relationship between query and document representations. Building upon the hypernetwork literature in machine learning (Ha et al., 2017; Sendera et al., 2023; von Oswald et al., 2020), we propose Hypencoder–a generic framework that learns a query-dependent multi-layer neural network as a similarity function that is applied to the document representations. In more detail, Hypencoder applies attention-based hypernetwork layers, called hyperhead layers, to the contextualized query embeddings output by a backbone transformer encoder. Each hyperhead layer produces the weight and bias matrices for a neural network layer in the query-dependent similarity network, called the q-net. The q-net is then applied to each document representation, which results in a scalar relevance score. We demonstrate that the Hypencoder framework can be optimized end-to-end and can be used for efficient retrieval from a large corpus. Specifically, we propose a graph-based greedy search algorithm that approximates exhaustive retrieval using Hypencoder while being substantially more efficient.
@@
We conduct extensive experiments on a wide range of datasets to demonstrate the efficacy of Hypencoder. We demonstrate that our implementation of Hypencoder for single vector document representations outperforms competitive single vector dense and sparse retrieval models on MS MARCO (Nguyen et al., 2016) and TREC Deep Learning Track data (Craswell et al., 2020b, a), in addition to complex retrieval tasks, such as TREC DL-Hard (Mackie et al., 2021), TREC Tip-of-the-Tongue (TOT) Track (Arguello et al., 2023), and the instruction following dataset FollowIR (Weller et al., 2024a). Across these benchmarks Hypencoder demonstrates consistent performance gain across experiments. Note that using the proposed approximation approach, retrieval from MS MARCO (Nguyen et al., 2016) with approximately 8.8 million documents only takes an average of 59.6 milliseconds per query on a single NVIDIA L40S GPU.
@@
A main advantage of hypernetworks in machine learning is their ability to learn generalizable representations. To demonstrate that Hypencoder also inherits this generalization quality, we evaluate our model under various domain adaptation settings: (1) adaptation to question answering datasets in biomedical and financial domains, and (2) adaptation to other retrieval tasks, including entity and argument retrieval, where Hypencoder again demonstrates superior performance compared to the baselines.
@@
We believe that these performance gains are just the by-product of our main contributions; Hypencoder introduces a new way to think about what retrieval and relevance functions can be, it opens a new world of possibilities by bridging the gap between neural networks and retrieval similarity functions. We believe Hypencoder is especially important at this time given the new demands for longer and more complex queries brought on by the widespread usage of large language models and it is our belief that Hypencoder represents an important step towards this goal. To help facilitate this goal we will open source all our code for training, retrieval, and evaluation. [^1] Available at https://github.com/jfkback/hypencoder-paper
@@ 2. Related Work
Vector-based models that use sparse vectors have existed for decades, with each index representing a term in the corpus vocabulary. Document-query similarity is computed using measures like $l_{2}$ distance, inner product, or cosine similarity, with term weighting methods such as TF-IDF being a substantial focus to improve performance. With the emergence of deep neural networks, focus shifted to learning representations for queries and documents. SNRM by Zamani et al. (2018) was the first deep learning model to retrieve documents from large corpora by learning latent sparse vectors. Following works leveraged pretrained transformer models like BERT (Devlin et al., 2019) using single dense vector representations (Karpukhin et al., 2020). Recent improvements have focused on training techniques including self-negative mining (Xiong et al., 2021; Qu et al., 2021; Prakash et al., 2021; Zhan et al., 2021), data augmentation (Qu et al., 2021; Lin et al., 2023a), distillation (Hofstätter et al., 2020; Lin et al., 2023a, 2020), corpus-pretraining (Gao and Callan, 2022; Izacard et al., 2022; Xiao et al., 2022; Ma et al., 2021), negative-batch construction (Hofstätter et al., 2021) and curriculum learning (Zeng et al., 2022; Lin et al., 2023b). Alternative approaches include ColBERT (Khattab and Zaharia, 2020), which uses multiple dense vectors, and SPLADE (Formal et al., 2021), which revisits sparse representations using pretrained masked language models.
Though these methods vary substantially, they all share a fundamental commonality, that relevance is based on an inner product (or in some cases cosine similarity). We believe that this is a significant limitation of these methods and one which hampers the performance of these models on complex retrieval tasks. Our method circumvents this limitation by learning a query-dependent small neural network that is fast enough to run on the entire collection (or used in an approximate way; see Section 3.6 for details).
@@
Light-weight relevance models using neural networks have demonstrated improved retrieval performance compared to simple methods like inner products. Early iterations came in the form of learning-to-rank models (Burges et al., 2005, 2006) which use query and document features to produce relevance scores for reranking. While these models traditionally used engineered features, more recent approaches adopted richer inputs. For instance, MatchPyramid (Pang et al., 2016) and KNRM (Xiong et al., 2017) use similarity matrices between non-contextualized word embeddings, while Duet (Mitra et al., 2017; Mitra and Craswell, 2019) combines sparse and dense term features in a multi-layer perceptron. DRMM (Yang et al., 2018) utilized histogram features as input to neural networks for scoring. Since the advent of BERT (Devlin et al., 2019), focus has shifted to making transformer models more efficient, such as PreTTR (MacAvaney et al., 2020) which separately precomputes query and document hidden states. Recently, LITE (Ji et al., 2024) extended ColBERT’s similarity using column-wise and row-wise linear layers for scoring.
In the recommender systems community, learned similarity measures have been widely used (He et al., 2017; He and Chua, 2017). The common usage of neural scoring methods in recommendation has inspired research into efficient retrieval with more learned scoring signals. For instance, BFSG (Tan et al., 2020) supports efficient retrieval with arbitrary relevance functions by using a graph of item representations and a greedy search strategy over nodes of the graph. A recent improvement on BFSG uses the scoring models gradient to prune directions that are unlikely to have relevant items (Zhao et al., 2024). Other works make use of queries to form a query-item graph to produce more informative neighbors (Tan et al., 2021).
Our work differs from these works in one major way, we do not have a query representation and document representation thus our method requires no combination step, instead we produce a query-conditioned neural network for each query and directly apply this to the document representation. This approach can reduce the similarity network’s size and does not require choosing between inference speed and larger query representations. Furthermore the flexibility of our framework means we can replicate any existing learned relevance model as discussed in Section 3.7. On a broader note there has been surprisingly little work on neural based scoring for full-scale retrieval, especially in the modern era of transformer based encoders. We hope our work can be a useful foundation and proof-of-concept for future work in this area.
@@ 3. Hypencoder
Neural ranking models can be generally categorized into early-interaction and late-interaction models (Dehghani et al., 2017; Nogueira et al., 2019; Khattab and Zaharia, 2020; Wang et al., 2023; Li et al., 2022; Gao et al., 2021; Hofstätter et al., 2022). Currently, the most common implementation of early-interaction models is in the form of cross-encoders (Figure 1 (second from the left)), where the query text $q$ and document text $d$ are concatenated (together with some predefined tokens or templates) and fed to a transformer network that learns a joint representation of query and document and finally produces a relevance score. The joint representation prevents these models from being able to precompute document representations, thus they cannot be used efficiently on large corpora (Guo et al., 2020; Mitra and Craswell, 2018).
@@
In order to take advantage of efficient indexing techniques, such as an inverted index in the case of sparse representations (Zamani et al., 2018; Formal et al., 2021) or approximate nearest neighbor (ANN) search in the case of dense representations (Karpukhin et al., 2020), many existing works use pooling techniques to obtain a single vector representation for each query and document and then employs simple and light-weight scoring functions, such as inner product or cosine similarity. There also exist more expensive methods that do not use pooling and perform such light-weight scoring functions at the vector level and then aggregate them, such as the maximum inner product similarity used in ColBERT (Khattab and Zaharia, 2020).
On the Limitations of Linear Similarity Functions (e.g., Inner Product)
@@
Let $C$ denote a corpus of $N$ documents, each being represented by an $h$-dimensional vector. A perfect ranking of documents in $C$ for a provided query is a ranking where all relevant documents are ranked above all non-relevant documents. According to Radon’s Theorem (Radon, 1921), any set of $h+1$ document vectors with $h$ dimensions can be partitioned into two sets whose convex hulls intersect. An important application of Radon’s Theorem is in calculating the Vapnik–Chervonenkis (VC) dimension (Vapnik and Chervonenkis, 1971) of $h$-dimensional vectors with respect to linear separability. For any $h+2$ vectors, the two subsets of a Radon partition cannot be linearly separated. In other words, for $N>h+1$, there exists at least one group of documents that is not linearly separable from the rest. In the real world, since $N\gg h+1$, there are indeed many such non-separable subsets. If any two of these subsets contain all the relevant documents for a query, then no linear similarity function can perfectly separate relevant from irrelevant documents. This includes inner product similarity and guarantees that, for some query, there will be an imperfect ranking.
@@ 3.1. Hypencoder Overview
An overview of our model is depicted in Figure 1 (right); it represents a new category of models that sit between a cross-encoder and a bi-encoder model. Like a bi-encoder model, our method computes the query and document representations separately, but unlike most existing retrieval methods, our method allows for more complicated matching signals like those present in cross-encoder models. Following existing methods, we have a query encoder and a document encoder. When a document $d$ is input into to the document encoder, we obtain a representation similar to existing encoder models, namely a set of one or more vectors $E_{d}\in\mathbb{R}^{m\times h}$ that represent the document’s content, where $m$ is the number of vectors and $h$ is the dimension of the vectors. Though we focus on vectors in this work, in theory, the representation can be anything a neural network can output.
Now comes our unique contribution that allows our method to consider more complex similarity signals. Given the query $q$, the query encoder $\Phi$ first produces a set of contextualized embeddings in a similar way to existing encoder models which we will call $E_{q}\in\mathbb{R}^{n\times h}$, where $n$ is the number of embeddings and $h$ is the dimension of the embeddings. At this point while existing methods apply a simple pooling mechanism, our query encoder instead uses a hyper-head. The hyper-head takes $E_{q}$ and produces a set of matrices and vectors that are then used as the weights and biases for a small neural network which we coin the q-net. The q-net is a query-dependent function for estimating relevance scores for each document, meaning each q-net is unique to the query that created it, unlike existing neural scoring methods which use a shared set of weights for all queries. To find the relevance of a document, the document representation $E_{d}$ is passed as input to the q-net which outputs the relevance score.
Hypencoder is a generic framework which allows direct application of existing paradigms from neural retrieval and, more broadly, machine learning. For example, Hypencoder could easily work with multiple vectors similar to existing multi-vector models, e.g., (Khattab and Zaharia, 2020), or use training routines popularized in dense retrieval, e.g., (Xiong et al., 2021; Prakash et al., 2021; Zeng et al., 2022; Lin et al., 2023a). As an initial exploration, this paper focuses on showing the efficacy of Hypencoder without additional complexity and thus uses a single vector document representation and no complex training recipes.

### v2

@@ Abstract.
Existing information retrieval systems are largely constrained by their reliance on vector inner products to assess query-document relevance, which naturally limits the expressiveness of the relevance score they can produce. We propose a new paradigm; instead of representing a query as a vector, we use a small neural network that acts as a learned query-specific relevance function. This small neural network takes a document representation as input (in this work we use a single vector) and produces a scalar relevance score. To produce the small neural network we use a hypernetwork, a network that produces the weights of other networks, as our query encoder. We name this category of encoder models Hypencoders. Experiments on in-domain search tasks show that Hypencoders significantly outperform strong dense retrieval models and even surpass reranking models and retrieval models with an order of magnitude more parameters. To assess the extent of Hypencoders’ capabilities, we evaluate on a set of hard retrieval tasks including tip-of-the-tongue and instruction-following retrieval tasks. On harder tasks, we find that the performance gap widens substantially compared to standard retrieval tasks. Furthermore, to demonstrate the practicality of our method, we implement an approximate search algorithm and show that our model is able to retrieve from a corpus of 8.8M documents in under 60 milliseconds.
retrieval models; neural ranking models; learning to rank
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729983isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrievalccs: Computing methodologies Machine learning
@@ 1. Introduction
Efficient neural retrieval models are based on a bi-encoder (or two-tower) architecture, in which queries and documents are represented separately using either high-dimensional sparse (Zamani et al., 2018; Formal et al., 2021, 2022) or relatively low-dimensional dense vectors (Karpukhin et al., 2020; Xiong et al., 2021; Khattab and Zaharia, 2020; Zeng et al., 2022; Gao and Callan, 2022). These models use simple and lightweight similarity functions, e.g., inner product or cosine similarity, to compute the relevance score for a given pair of query and document representations. We demonstrate theoretically that inner product similarity functions fundamentally limit the types of relevance that retrieval models can express. Specifically, we prove that there is always a set of relevant documents which cannot be perfectly retrieved regardless of the query vector and specific encoder model.
@@
Motivated by this theoretical argument, we introduce a new category of retrieval models that can capture complex relationships between query and document representations. Building upon the hypernetwork literature in machine learning (Ha et al., 2017; Sendera et al., 2023; von Oswald et al., 2020), we propose Hypencoder–a generic framework that learns a query-dependent multi-layer neural network as a similarity function that is applied to the document representations. In more detail, Hypencoder applies attention-based hypernetwork layers, called hyperhead layers, to the contextualized query embeddings output by a backbone transformer encoder. Each hyperhead layer produces the weight and bias matrices for a neural network layer in the query-dependent similarity network, called the $q$-net. The $q$-net is then applied to each document representation, which results in a scalar relevance score. We demonstrate that the Hypencoder framework can be optimized end-to-end and used for efficient retrieval from a large corpus. Specifically, we propose a graph-based greedy search algorithm that approximates exhaustive retrieval using Hypencoder while being substantially more efficient.
@@
We conduct extensive experiments on a wide range of datasets to demonstrate the efficacy of Hypencoder. We show that our implementation of Hypencoder outperforms strong single-vector dense and sparse retrieval models on MSMARCO (Nguyen et al., 2016) and TREC Deep Learning Track data (Craswell et al., 2020b, a), in addition to complex retrieval tasks, such as TREC DL-Hard (Mackie et al., 2021), TREC Tip-of-the-Tongue (TOT) Track (Arguello et al., 2023), and the instruction following dataset FollowIR (Weller et al., 2024a). Across these benchmarks, Hypencoder shows consistent performance gains. Using the proposed approximate retrieval approach on MSMARCO (Nguyen et al., 2016), with approximately 8.8 million passages, results in an average search latency of 59.6 milliseconds per query on a single NVIDIA L40S GPU.
@@
A main advantage of hypernetworks in machine learning is their ability to learn generalizable representations. To demonstrate that Hypencoder also inherits this generalization quality, we evaluated our model under various domain adaptation settings: (1) adaptation to question answering datasets in biomedical and financial domains, and (2) adaptation to other retrieval tasks, including entity and argument retrieval, where Hypencoder again demonstrates superior performance compared to the baselines.
@@
We believe that these performance gains are a byproduct of our main contributions: Hypencoder introduces a new way to think about what retrieval and relevance functions can be, it opens a new world of possibilities by bridging the gap between neural networks and retrieval similarity functions. We believe Hypencoder is especially important at this time given the new demands for longer and more complex queries brought on by the widespread usage of large language models and it is our belief that Hypencoder represents an important step towards this goal. To help facilitate this goal, we have open sourced all of our code for training, retrieval, and evaluation. [^1] Available at https://github.com/jfkback/hypencoder-paper
@@ 2. Related Work
Vector-based models that use sparse vectors have existed for decades, with each index representing a term in the corpus vocabulary. Document-query similarity is computed using measures like $\ell^{2}$ distance, inner product, or cosine similarity, with term weighting methods such as TF-IDF being a substantial focus to improve performance. With the emergence of deep neural networks, focus shifted to learning representations for queries and documents. SNRM by Zamani et al. (2018) was the first deep learning model to retrieve documents from large corpora by learning latent sparse vectors. Following works leveraged pretrained transformer models like BERT (Devlin et al., 2019) using single dense vector representations (Karpukhin et al., 2020). Recent improvements have focused on training techniques including self-negative mining (Xiong et al., 2021; Qu et al., 2021; Prakash et al., 2021; Zhan et al., 2021; Zeng et al., 2023), data augmentation (Qu et al., 2021; Lin et al., 2023a), distillation (Hofstätter et al., 2020; Lin et al., 2023a, 2020; Zeng et al., 2025), corpus-pretraining (Gao and Callan, 2022; Izacard et al., 2022; Xiao et al., 2022; Ma et al., 2021), negative-batch construction (Hofstätter et al., 2021) and curriculum learning (Zeng et al., 2022; Lin et al., 2023b). Alternative approaches include ColBERT (Khattab and Zaharia, 2020), which uses multiple dense vectors, and SPLADE (Formal et al., 2021), which revisits sparse representations using pretrained masked language models.
Though these methods vary substantially, they all share a fundamental commonality, that relevance is based on an inner product (or in some cases cosine similarity). We believe that this is a significant limitation of these methods and one which hampers the performance of these models on complex retrieval tasks. Our method circumvents this limitation by learning a small query-dependent neural network that is fast enough to run on the entire collection (or approximately run; see Section 3.6 for details).
@@
Lightweight relevance models using neural networks have demonstrated improved retrieval performance compared to simple methods like inner products. Early iterations came in the form of learning-to-rank models (Burges et al., 2005, 2006) which use query and document features to produce relevance scores for reranking. While these models traditionally used engineered features, more recent approaches adopted richer inputs. For instance, MatchPyramid (Pang et al., 2016) and KNRM (Xiong et al., 2017) use similarity matrices between non-contextualized word embeddings, while Duet (Mitra et al., 2017; Mitra and Craswell, 2019) combines sparse and dense term features in a multi-layer perceptron. DRMM (Yang et al., 2018) utilized histogram features as input to neural networks for scoring. Since the advent of BERT (Devlin et al., 2019), focus has shifted to making transformer models more efficient, such as PreTTR (MacAvaney et al., 2020) which separately precomputes query and document hidden states. Recently, LITE (Ji et al., 2024) extended ColBERT’s similarity using column-wise and row-wise linear layers for scoring.
In the recommender systems community, learned similarity measures have been widely used (He et al., 2017; He and Chua, 2017). The common usage of neural scoring methods in recommendation has inspired research into efficient retrieval with more learned scoring signals. For instance, BFSG (Tan et al., 2020) supports efficient retrieval with arbitrary relevance functions by using a graph of item representations and a greedy search strategy. A recent improvement on BFSG uses the scoring models gradient to prune directions that are unlikely to have relevant items (Zhao et al., 2024). Other works make use of queries to form a query-item graph to produce more informative neighbors (Tan et al., 2021).
Our work differs from these works in one major way, the weights of neural scoring model are dynamically produced for each query. This means the full capacity of the scoring model can be dedicated to just the features which are meaningful for that specific query and that no additional combination step is needed to aggregate the query and document representations as the query representation is already encoded in the scoring model. Furthermore, the flexibility of our framework means we can replicate any existing learned relevance model as discussed in Section 3.7. On a broader note, there has been surprisingly little work on neural-based scoring for full-scale retrieval, especially in the transformer era. We hope our work can be a useful foundation and proof-of-concept for future work in this area.
@@ 3. Hypencoder
Neural ranking models can be generally categorized into early-interaction and late-interaction models (Dehghani et al., 2017; Nogueira et al., 2019; Khattab and Zaharia, 2020; Wang et al., 2023; Li et al., 2022; Gao et al., 2021; Hofstätter et al., 2022). Currently, the most common implementation of early-interaction models is in the form of cross-encoders (Figure 1 (second from the left)), where the query text $q$ and document text $d$ are concatenated (together with some predefined tokens or templates) and fed to a transformer network that learns a joint representation of query and document and finally produces a relevance score. The joint representation prevents these models from being able to precompute document representations, so they cannot be used efficiently on large corpora (Guo et al., 2020; Mitra and Craswell, 2018).
@@
In order to take advantage of efficient indexing techniques, such as an inverted index or approximate nearest neighbor (ANN) search, many existing works use pooling techniques to obtain a single vector representation for each query and document and then employ simple and lightweight scoring functions, such as inner product or cosine similarity. There also exist more expensive methods that do not use pooling and perform such lightweight scoring functions at the vector level and then aggregate them, such as the maximum inner product similarity used in ColBERT (Khattab and Zaharia, 2020).
On the Limitations of Linear Similarity Functions such as Inner Product
@@
Let $C$ denote a corpus of $N$ documents, each represented by an $h$-dimensional vector. A perfect ranking of documents in $C$ for a provided query is a ranking where all relevant documents are ranked above all non-relevant documents. According to Radon’s Theorem (Radon, 1921), any set of $h+2$ document vectors of dimension $h$ can be partitioned into two sets whose convex hulls intersect. An important application of Radon’s Theorem is in calculating the Vapnik–Chervonenkis (VC) dimension (Vapnik and Chervonenkis, 1971) of $h$-dimensional vectors with respect to linear separability. For any $h+2$ vectors, the two subsets of a Radon partition cannot be linearly separated. In other words, for $N>h+1$, there exists at least one group of documents that is not linearly separable from the rest. In the real world, since $N\gg h+1$, there are indeed many such non-separable subsets. If any two of these subsets contain all the relevant documents for a query, then no linear similarity function can perfectly separate relevant from irrelevant documents. This includes inner product similarity and guarantees, that for some query, there will be an imperfect ranking.
@@ 3.1. Hypencoder Overview
An overview of our model is depicted in Figure 1 (right); it represents a new category of models that sit between a cross-encoder and a bi-encoder model. Like a bi-encoder model, our method computes the query and document representations separately, but unlike most existing retrieval methods, our method allows for more complicated matching signals like those present in cross-encoder models. Following existing methods, we have a query encoder and a document encoder. When a document $d$ is input into the document encoder, we obtain a representation similar to existing encoder models, namely a set of one or more vectors $E_{d}\in\mathbb{R}^{m\times h}$ that represent the document’s content, where $m$ is the number of vectors and $h$ is the dimension of the vectors. Although we focus on vectors in this work, in theory, the representation can be anything a neural network can output.
Now comes our unique contribution, which allows our method to consider more complex similarity signals. Given the query $q$, the query encoder $\Phi$ first produces a set of contextualized embeddings in a similar way to existing encoder models which we will call $E_{q}\in\mathbb{R}^{n\times h}$, where $n$ is the number of embeddings and $h$ is the dimension of the embeddings. At this point, existing methods apply a simple pooling mechanism; in contrast, our query encoder uses a hyper-head. The hyper-head takes $E_{q}$ and produces a set of matrices and vectors that are then used as the weights and biases for a small neural network which we coin the $q$-net. The $q$-net is a query-dependent function to estimate relevance scores for a document, meaning each $q$-net is unique to the query that created it, unlike existing neural scoring methods, which use a shared set of weights for all queries. To find the relevance of a document, the document representation $E_{d}$ is passed as input to the $q$-net which outputs the relevance score.
Hypencoder is a generic framework that allows direct application of existing paradigms from neural retrieval and, more broadly, machine learning. For example, Hypencoder could easily work with multiple vectors similar to existing multi-vector models (Khattab and Zaharia, 2020), or use training routines popularized in dense retrieval, e.g., (Xiong et al., 2021; Prakash et al., 2021; Zeng et al., 2022; Lin et al., 2023a). As an initial exploration, this paper focuses on showing the efficacy of Hypencoder without additional complexity, and thus uses a single vector document representation and no complex training recipes.